A. Discuss any key considerations that you may pay attention to when making



investment decisions for financial management.




B. Hankede is interested in determining the time period for his K3,600 to accumulate



to K4,000 if saved today. How long will it take for Hankede to achieve his objective



with semiannual compound rate of interest of 6% per annum?




C. A company has borrowed K800,000 from a bank. The loan is to be repaid by level



instalments, payable annually in arrears for 10 years from the date the loan is



made. The annual repayments are based on an effective rate of interest of 8% per



annum.



I. Calculate the amount of the level annual payment which will be paid over



the 10-year term.



II. Construct an amortization schedule showing the capital and interest



components over the first four years of the loan.

1
Expert's answer
2022-02-08T12:29:03-0500

1.

a). Rate of inflation- this is important to the investor since he will be able to know if the invested money will increase in value or decrease in value. If there is inflation, then the investor should not risk his money.

b). Risks involved- if the risks present are of high level, then the investor should not risk investing his money there is a chance they will be lost.

c). Return on investment- this refers to amount that the investor will get back after sometime. If the business is likely to give more returns then the investor should invest.

d). Investment period- this refers to the period in which the investor will get his returns back. The shorter the period, the better the venture.


2. Interest "A=p(1+r\/n)"


"4000=3600(1+0.06\/n)"

"= 0.5 years"

=6 months


3. "I=p*r*t"

"=800,000\u00d70.08\u00d710"

"=1,440,000"

"= 1,440,000\u00f710"

"= 144000" per year.